OpenStudy (anonymous):

A culture of bacteria is growing at a rate of 3e ^{.2t} per hour, with t in hours and 0<=t<=20. How many new bacteria will be in the culture after the first five hours?

OpenStudy (anonymous):

I have\[ Q'(t) =3e ^{.2t}\] \[Q(t) = 15e ^{.2t} + c\]

OpenStudy (anonymous):

is that correct thinking?

OpenStudy (anonymous):

if you're given the growth rate, then use the integral to find the amount at time t i do not believe you integrate correctly

OpenStudy (anonymous):

integrate from 0 to t where t =5

OpenStudy (anonymous):

ohhh yeah that would make sense

OpenStudy (anonymous):

hmm ok so that comes out with the correct answer of 25, but when I try it for the next part b) How many new bacteria are introduced in the sixth through the fourteenth hours? I come up with the answer 196 when it should be 205

OpenStudy (anonymous):

integrate from 6 to 14 either you integrated incorrectly or the book is wrong usually its a calculation error so double check that

OpenStudy (anonymous):

what happens to the + c on the integration? Does it mean anything?

OpenStudy (anonymous):

\[Q(0) = 0 = 15e ^{.2t} + c\] so shouldn't c = -15 and be used when calculating?

OpenStudy (anonymous):

i said your integration was incorrect where in teh world did you get 15

OpenStudy (anonymous):

3/.2 = 15?

OpenStudy (anonymous):

nvm i misread the .2 as a 2 my mistake

OpenStudy (anonymous):

unless Q(0)=0 was given to you, or any other values, you cant really find c

OpenStudy (anonymous):

well I Q(t) represents the new bacteria so I assumed that Q(0) would = 0 right?

OpenStudy (anonymous):

well no, new bacteria means the c would cancel out, so essentially c=0

OpenStudy (anonymous):

well then at Q(0) we would already have 15 bacteria, that doesn't make sense

OpenStudy (anonymous):

new bacteria, meaning, over a period of time

OpenStudy (anonymous):

you are not calculating for a single point in time

OpenStudy (anonymous):

after hte first 5 hours means \[15e^{.2*5}+c-(15e^{.2*0}+c)\] i think

OpenStudy (anonymous):

ok that makes sense, I guess I'm just a little shaky on my understanding of the constant of integration is all.

OpenStudy (anonymous):

so for the second part we have \[15e ^{.2 * 14} - (15e ^{.2 * 6}) = 246.67 - 49.80 = 196.87?\]